40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 103     Entry time: Tue Nov 25 17:30:32 2008
Author: Aidan 
Type: Laser 
Category: Fiber 
Subject: Preliminary noise budgets for fibre stabilization 

Here are some preliminary noise budgets for the FS experiment ... or rather, the setup shown in attached figure. Looks like the laser frequency noise dominates.
Not surprising given that the optical path length difference is ~150m.

The photodiode noise is not particularly high, although if we stabilize the frequency and intensity of the NPRO it might become a factor.

Anyway, the phase and frequency noise plots are attached. Will post references to the origin of the actual curves for the different noise sources soon.
Attachment 1: simulation_v1B.jpg  70 kB  | Hide | Hide all
simulation_v1B.jpg
Attachment 2: phaseNoiseBudget-NPRO-PDA255.pdf  16 kB  | Hide | Hide all
phaseNoiseBudget-NPRO-PDA255.pdf
Attachment 3: freqNoiseBudget-NPRO-PDA255.pdf  22 kB  | Hide | Hide all
freqNoiseBudget-NPRO-PDA255.pdf
ELOG V3.1.3-